OpenStudy (anonymous):

Write an equation in standard form of the line that passes through the given pion and has the given slope m. Ex: (4,3),m=7 using the formula ax+by=c

OpenStudy (lncognlto):

y = mx + c ... (1) Now, as m = 7, x = 4, and y = 3, we substitute these values in (1). Therefore, 3 = (7)(4) + c Therefore, c = 3 - 28 Therefore, c = -25 Therefore y = 7x -25 or y - 7x = -25.